OpenStudy (mopsy510):

Can you solve this? The radius of a circle is 20 kilometers. What is the angle measure of an arc 5π kilometers long? @ranga

OpenStudy (anonymous):

Length of an arc = n/360 (2pi)(radius) 5 pi = n/360 (2pi)(20) 5 pi = n/360 (40 pi) 1/8 = n/360 n = 45 so 45 degrees.

OpenStudy (mopsy510):

How about this one: The radius of a circle is 2 centimeters. What is the length of a 45° arc?

OpenStudy (anonymous):

length of the arc = n/360 (2 pi)(radius) length of the arc = 45/360 (2pi)(2) Length of the arc = 1/8 (pi) = pi/8

OpenStudy (anonymous):

45/360 = 1/8

OpenStudy (anonymous):

Your welcome.

OpenStudy (mopsy510):

The diameter of a circle is 40 feet. What is the angle measure of an arc bounding a sector with area 37π square feet?

OpenStudy (mopsy510):

@Easyaspi314

OpenStudy (anonymous):

I am working with students that have the courtesy to say "thank you" after spending time helping them.

OpenStudy (anonymous):

Good luck!